OpenStudy (anonymous):

-x/3+5>8 this is rally hard for me

OpenStudy (anonymous):

help me kevin please, ill do anything I gotta get outta summer school!

OpenStudy (anonymous):

don't worry, it's comming :)

OpenStudy (anonymous):

I got something thats really hard for you

OpenStudy (anonymous):

What is 2x2*3+7x-4x=9?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

that sounded rapey

OpenStudy (anonymous):

what kevin had said?

OpenStudy (anonymous):

\[-\frac{x}{3} + 5 > 8\] \[~~~~~-\frac{x}{3} > 3\] \[~~~~~~-x > 9\] \[~~~~~~~~~x < -9\]

OpenStudy (anonymous):

everything is rape

OpenStudy (anonymous):

kevin is my boyfriend name so it feels right

OpenStudy (anonymous):

x<-9 (:

OpenStudy (anonymous):

ITT people who love rape

OpenStudy (anonymous):

I got something that would make you feel right

OpenStudy (anonymous):

Sorry, I'm slow at typing

OpenStudy (anonymous):

its ok hun! ;)

OpenStudy (anonymous):

X<-9

OpenStudy (anonymous):

rofl

OpenStudy (anonymous):

keonia you won u get a medal

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

ok I got you a medal, when people get my questions right they get a medal so look out im about to post another!

OpenStudy (anonymous):

Scroll up some. I think you missed mine. I don't care about the medals, but I think you should take a look at the work.

OpenStudy (anonymous):

o.o kevin is dat you.....

OpenStudy (anonymous):

who is single

OpenStudy (anonymous):

im a thirsty female

OpenStudy (anonymous):

the thirst is real

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wt...

OpenStudy (anonymous):

i was saying thank you for the medal